### v1

# Improbable Bigrams Expose Vulnerabilities of Incomplete Tokens in Byte-Level Tokenizers

Abstract
Tokenization is a crucial step that bridges human-readable text with model-readable discrete tokens. However, recent studies have revealed that tokenizers can be exploited to elicit unwanted model behaviors. In this work, we investigate incomplete tokens, i.e., undecodable tokens with stray bytes resulting from byte-level byte-pair encoding (BPE) tokenization. We hypothesize that such tokens are heavily reliant on their adjacent tokens and are fragile when paired with unfamiliar tokens. To demonstrate this vulnerability, we introduce improbable bigrams: out-of-distribution combinations of incomplete tokens designed to exploit their dependency. Our experiments show that improbable bigrams are significantly prone to hallucinatory behaviors. Surprisingly, alternative tokenizations of the same phrases result in drastically lower rates of hallucination (93% reduction in Llama3.1). We caution against the potential vulnerabilities introduced by byte-level BPE tokenizers, which may impede the development of trustworthy language models.
Improbable Bigrams Expose Vulnerabilities of Incomplete Tokens in Byte-Level Tokenizers
Eugene Jang1     Kimin Lee2     Jin-Woo Chung1     Keuntae Park1     Seungwon Shin2 1S2W Inc.    2KAIST 1{genesith,jwchung,keuntae.park}@s2w.inc    2{kiminlee,claude}@kaist.ac.kr
1 Introduction
Tokenization is an important step in the large language model (LLM) pipeline, serving as the bridge between text inputs and the discrete tokens processed by the model. Improper tokenization can lead to undesirable behaviors, such as glitch token hallucinations Rumbelow and Watkins (2023); Land and Bartolo (2024) and errors in numerical reasoning Singh and Strouse (2024). Furthermore, it has been observed that tokenizers can introduce elements of unfairness Petrov et al. (2023) and bias Ovalle et al. (2024) into models. Given the increasing duration and cost of model training, investigating and understanding the impact of the tokenizer is becoming increasingly important.
Figure 1: An improbable bigram phrase that combines two incomplete tokens to cause hallucinatory behaviors in the Qwen2.5 model. This behavior persists across multiple models and with well-trained tokens. An alternative tokenization of the same phrase does not cause hallucinations.
Recent works specifically analyze the tokenization step to identify inputs that provoke unwanted model behaviors. Land and Bartolo (2024) propose embedding layer heuristics to identify glitch tokens — undertrained tokens that cause hallucinations and also enable jailbreaks Geiping et al. (2024). Wang et al. (2024) create adversarial questions designed to induce incorrect segmentation by the tokenizer, which degrades model performance.
In this work, we investigate a specific vulnerability associated with incomplete tokens in byte-level byte pair encoding (BPE) tokenizers. These tokens, also known as undecodable tokens, are byte-level tokens that cannot be decoded independently and must appear in conjunction with certain other tokens to form legal Unicode characters. We explore the fragility of these tokens by constructing improbable bigrams, which are unlikely yet permissible combinations of two incomplete tokens. Similar to glitch tokens, improbable bigrams cause hallucinations to benign user requests, as seen in Figure 1. However, unlike glitch tokens, improbable bigrams can cause hallucinatory behavior even when their constituent tokens are well-trained.
Our experiments across multiple LLM families (such as Llama-3.1, Qwen2.5, and Mistral-Nemo) demonstrate that these bigrams are significantly more prone to producing hallucinations compared to bigrams formed from complete tokens. Intriguingly, alternative tokenizations of the same phrases rarely exhibit hallucinations. We believe that our findings contribute to the growing body of research on understanding the potential weaknesses of BPE tokenization, with implications for the development of more robust language models.
2 BPE Tokenization
Tokenizer Vocab Size Incomplete Tokens Incomplete Bigrams
Meta-Llama-3.1 128k 1224 71k
Exaone-3.0 102k 1222 36k
Qwen2.5 151k 1320 39k
Mistral-Nemo 131k 1307 135k
Command-R-v01 255k 2956 1479k
Table 1: Vocabulary sizes and number of incomplete tokens in different tokenizers. Incomplete bigrams refer to the total number of legal bigrams that can be created by combining two incomplete tokens.
Byte pair encoding (BPE), originally developed as a data compression algorithm Gage (1994), has evolved into a popular tokenization scheme used by many modern language models Sennrich et al. (2016). BPE tokenization iteratively expands the vocabulary by selecting frequent pairs of existing tokens as new vocabulary items. In particular, byte-level BPE, which applies the BPE algorithm at the byte-level rather than the character level Wang et al. (2019), offers distinct advantages. Firstly, byte-level tokenization eliminates out-of-vocabulary issues by representing all Unicode characters (154,998 in Unicode v16.0) as combinations of 256 base bytes, ensuring comprehensive coverage. Secondly, it allows for more efficient data compression. The widespread adoption of byte-level tokenization in cutting-edge models such as GPT-4 and Llama 3.1 underscores its effectiveness.
While some have attributed BPE’s effectiveness to its compression Gallé (2019); Goldman et al. (2024), recent work has challenged this assumption Schmidt et al. (2024). Bostrom and Durrett (2020) point out that greedy compression prioritizes frequency over linguistically meaningful boundaries. This observation is supported by other works that report model improvements from morphologically motivated modifications to BPE tokenizers Limisiewicz et al. (2024); Bauwens and Delobelle (2024). These works demonstrate that incorrect segmentation by BPE tokenization cause suboptimal model performance.
3 Incomplete Tokens and Improbable Bigrams
{CJK*}
UTF8gbsn
Byte-level BPE tokenizers are agnostic not only to morphemes but also to character boundaries. Training with the BPE algorithm on texts where characters are encoded with multiple bytes often results in the formation of incomplete tokens. Incomplete tokens contain stray bytes (like <0x9F> in the token ‘<0x9F>能’) that do not correspond to recognizable characters. These incomplete tokens constitute a small but significant portion of the byte-level vocabulary. Table 1 presents the number of discovered incomplete tokens in the vocabularies of various models employing byte-level tokenizers (see Appendix A for details).
Task User prompt
Definition What does ‘‘{Phrase}’’ mean?
Do You Know? Today I heard about ‘‘{Phrase}’’. Do you know what this means?
Usernames [Uncaptioned image]
Table 2: Prompt templates used to test phrase-level hallucinations of target phrases. At test time, {Phrase} is replaced with the tested phrase. If the model fails to repeat the phrase in all three prompts, we consider the phrase to be hallucinatory.
Figure 2: Incomplete tokens are analyzed for their structure to find potential bigram candidates. C-bytes (continuation bytes) are colored in blue. Tokens with compatible structures are tested for viability and reachability through an decode-encode check.
While incomplete tokens are indistinguishable from complete tokens from the model’s perspective during training, their structure can influence how they are trained in subtle ways. First, incomplete tokens can only occur alongside other incomplete or single-byte tokens, as the “stray bytes” of incomplete tokens require additional bytes to form a viable character. This restriction limits the range of tokens they can be associated with. Second, the potential characters that these stray bytes can form may have little to no semantic overlap, leading to a representation that is inherently ambiguous and heavily reliant on neighboring tokens for disambiguation. We hypothesize that these factors exacerbate one another, causing incomplete tokens to become overly dependent on their context and brittle in response to unfamiliar adjacent tokens.
To demonstrate this vulnerability, we analyze the structure of incomplete tokens, categorizing those that end with stray bytes as prefixes and those that begin with stray bytes as suffixes. We further label prefixes and suffixes by the number of stray bytes they require to form a complete character, to find potential bigram combinations. We create adversarial combinations of these tokens, termed improbable bigrams, as illustrated in Figure 2. These bigrams exploit the dependency of two incomplete tokens to complete each other. We note that not all combinations are viable due to tokenization rules and merge priorities. Table 1 lists the number of all possible legal bigrams created from incomplete tokens across different tokenizers. To identify bigrams that are particularly out-of-distribution, we use the heuristic of multilinguality: if a bigram combines to form scripts from different languages, we assume it is highly unlikely to have been encountered during training. In our experiments, we demonstrate that improbable bigrams are prone to hallucinations.
4 Hallucination Experiments
Models We experiment using five recently released instruction-trained models that use byte-level BPE tokenization: Meta-Llama-3.1-8B-Instruct [^1] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct, EXAONE-3.0.-7.8B-Instruct [^2] https://huggingface.co/LGAI-EXAONE/EXAONE-3.0-7.8B-Instruct, Qwen2.5-32B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-32B-Instruct, Mistral-Nemo-Instruct-2407444https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407, and C4AI-Command-R-v01555https://huggingface.co/CohereForAI/c4ai-command-r-v01. We use greedy decoding for all experiments.
Models Improbable Bigrams Baseline Bigrams Llama 3.1 43/100 (43%) 0/100 (0%) Exaone 79/100 (79%) 26/100 (26%) Qwen2.5 38/100 (38%) 2/100 (2%) Mistral-Nemo 52/71 (73%) 0/71 (0%) Command-R 56/100 (56%) 9/100 (9%)
Table 3: Number of hallucinations by improbable bigram phrases (with incomplete tokens) and by baseline bigram phrases (with complete tokens).
Evaluation Prompts We assess each model’s ability to accurately process prompts containing improbable bigram phrases. We use three specific prompt templates designed to induce the model to repeat a target phrase from the input, regardless of the phrase’s nonsensical or unusual nature, as detailed in Table 2. We consider a phrase to have induced hallucinations only if the model fails to correctly repeat it across all three prompt templates.
Improbable Bigrams For each model, we construct improbable bigrams using the procedure outlined in Section 3. Noting that undertrained tokens often lead to hallucinations, we focus our experiments on well-trained incomplete tokens to isolate other factors. Employing the embedding heuristic methodology from Land and Bartolo (2024), we rank all tokens by their degree of training and select only the incomplete tokens in the upper half of this ranking. We then create up to 100 improbable bigrams for each model for our experiments.
Baselines To better demonstrate the fragility of improbable bigrams, we create a baseline of bigrams constructed with complete tokens. For each improbable bigram, we create a complete token counterpart by replacing the prefix and suffix with similarly-trained tokens. That is, from an initial incomplete token that is the i-th token in the ranked vocabulary, we find a complete token that is slightly less trained (e.g. the (i-1)-th token). This method ensures that each improbable bigram is compared to a baseline of complete bigrams with a similar level of undertraining.
Results As shown in Table 3, improbable bigrams suffer from significantly higher rates of hallucination than their complete token counterparts across all tested models. This is distinct from hallucinations involving glitch tokens, as previously investigated by Land and Bartolo (2024), which was attributed to undertrained tokens. These results suggest that even well-trained incomplete tokens can struggle to faithfully represent textual inputs.
5 Dehallucinations via Alternative Tokenization
Models Incomplete Token Hallucinations Alternative Tokenization Hallucinations Llama 3.1 0.43 0.03 (↓93%) Exaone 0.79 0.52 (↓34%) Qwen2.5 0.38 0.14 (↓63%) Mistral-Nemo 0.73 0.01 (↓98%) Command-R 0.56 0.57
Table 4: Frequency of hallucinations from improbable bigram phrases with original tokenization (Incomplete Tokens) and presegmented tokenization (Alternative Tokenization). Reduction of hallucinations by alternative tokenization is denoted by ↓ when applicable.
While we have hypothesized that improbable bigram hallucinations are caused by incomplete tokens, another possible explanation for their hallucinations is that they are simply just out-of-distribution. To test this hypothesis, we repeat the experiments using the same phrases, but with a different tokenization method that avoids incomplete tokens. Instead of tokenizing the bigram phrase according to the tokenizer, we avoid character boundary-crossing tokenization using pre-segmentation. {CJK*}UTF8ipxm For instance, the improbable bigram phrase “サーミ能” is normally tokenized into incomplete tokens ‘サー<0xE3><0x83>’ and ‘<0x9F>能’. However, it can be presegmented to isolate the character formed from stray bytes (‘ミ’). We split the string into three parts (‘サー’, ‘ミ’, and ‘能’) and tokenize each separately. Afterwards, the tokens can be appended together to create an alternative token representation of the same phrase.
Compared to improbable bigrams, alternative tokenization sequences can be suboptimal in two important ways. First, these sequences cannot be naturally generated by the tokenizer, ensuring that they are out-of-distribution for the model. Second, the sequences consist of more than two tokens, requiring the model to recall more tokens correctly to repeat the phrase accurately. If the out-of-distribution hypothesis holds, we would expect these phrases to induce more hallucinations with this alternative tokenization.
Table 4 shows results of the experiment repeated using the alternative tokenization scheme. Surprisingly, we find that these token sequences, which are both unnatural and longer, hallucinate significantly less than the original tokenization for most models. These results support our hypothesis that the hallucinatory behaviors are being caused by the incomplete tokens.
6 Conclusion
Through improbable bigrams, we demonstrate vulnerabilities of incomplete tokens present in byte-level BPE tokenizers. We conclude that incomplete tokens are significantly more prone to hallucinatory behaviors compared to complete tokens. Our findings suggest that model developers must be mindful of potential unwanted behaviors caused by incomplete tokens.
Limitations
This work provides evidence to suggest the fragility of incomplete tokens. We assess phrase-level hallucinations as a proxy for larger potential harms of model misbehavior. We limit our investigations to phrase-level hallucinations, rather than factual hallucinations. While we occasionally observe that the models confidently assert incorrect explanations about what is likely non-existent terminology, this was not possible to evaluate systematically. One reason is that the tested phrases contain scripts spanning many languages beyond the expertise of the authors. Another is that the baseline levels of factual hallucination when asked about ordinary terms in low-resource languages are already very high. In pilot experiments, we found that even models with high conversational fluency often make mistakes in explaining non-existent terms.
Our findings on improbable bigrams causing hallucinations were consistent in all 5 models. However, our secondary experiment on alternate tokenization did not cause improvements for the Cohere Command-R model. We hypothesize that the training of the tokens may have been different due to the model’s significantly larger vocabulary size, which is approximately double the other studied models. However, a more detailed investigation to verify this is left to future work. There is a possibility that these findings may be influenced by other unknown dependencies relating to tokenizer parameters and training parameters.
Appendix A Incomplete Bigram Construction
UTF-8 characters have a flexible number of bytes. Each multi-byte character is comprised of a starting byte and continuation bytes. Starting bytes indicate how many bytes the character is composed of. For example, the starting byte of a 3-byte character implies the following two characters will be continuation bytes. A more detailed explanation of the UTF-8 protocol and byte structures can be found in Limisiewicz et al. (2024).
Using starting bytes and continuation bytes, we can identify whether a byte sequence requires more continuation bytes to be fulfilled or has excess continuation bytes. As illustrated in Figure 2, each incomplete token can be analyzed for their structure. We test all possible combinations of incomplete bigrams by pairing bigrams of complementary structures. We focus only on bigrams that users could use to induce the intended incomplete bigram structure. Note that the resulting character of conjoined bytes may not always correspond to a usable Unicode character. There is also a possibility that the resulting phrase is tokenized into a different sequence of tokens. We test this by performing a decode-encode test to ensure the resulting string is valid and indeed tokenizes to the intended prefix and suffix.